1 litre flask containing NH3 at 2.0 atm at 200 K is connected by a narj-ow tube of negligible volume to another 800 mL flask containing HCl(g) at 8.0 atm at 200 K. The two gases react according to equation NH3 + HCl (g) NHXI 4 (S) AH = - 43 kJ/mol If .heat capacity of HCl(g) Cv is 20 J K 1 mol ', determine the heat produced

Grade:11

1 Answers

Arun
3 years ago
NH 
3
​ 
 (g)+HCl(g)→NH 
4
​ 
 Cl(s)
Mole 0.08 0.26 0
     0 0.18 0.08
⇒Q=0.08×4300=3440J
Also, Q=nC 
v
​ 
 ΔT⇒ΔT= 
0.18×20
3440
​ 
 =955.55
final
​ 
 (T 
∗2
​ 
 )=1255.55K
final
​ 
 (P 
∗2
​ 
 )= 
1.8
0.18×0.082×1255.55
​ 
 
=10.3 atmosphere